# Is the Lecture Engaging for Learning? Lecture Voice Sentiment Analysis for Knowledge Graph-Supported Intelligent Lecturing Assistant (ILA) System

### v1

@@ Abstract
This paper introduces an intelligent lecturing assistant (ILA) system that utilizes a knowledge graph to represent course content and optimal pedagogical strategies. The system is designed to support instructors in enhancing student learning through real-time analysis of voice, content, and teaching methods. As an initial investigation, we present a case study on lecture voice sentiment analysis, in which we developed a training set comprising over 3,000 one-minute lecture voice clips. Each clip was manually labeled as either engaging or non-engaging. Utilizing this dataset, we constructed and evaluated several classification models based on a variety of features extracted from the voice clips. The results demonstrate promising performance, achieving an F1-score of 90% for boring lectures on an independent set of over 800 test voice clips. This case study lays the groundwork for the development of a more sophisticated model that will integrate content analysis and pedagogical practices. Our ultimate goal is to aid instructors in teaching more engagingly and effectively by leveraging modern artificial intelligence techniques.
Keywords AI in Teaching and Learning  $\cdot$ Lecturing Assistant  $\cdot$ Voice Sentiment Analysis  $\cdot$ Knowledge Graph for Teaching and Learning  $\cdot$ Intelligent Educational Systems
1 Introduction
Engaging students and facilitating the retention of knowledge in long-term memory are fundamental aspects of effective teaching, yet many lectures often fail to achieve this goal. Although extensive research in cognitive science [1] and neuroscience [2] has suggested a multitude of scientifically-based strategies for effective teaching, the application of these findings in real classrooms remains limited. Advancements in artificial intelligence offer a promising avenue for transforming the science of learning to real-world experience. This paper introduces a novel knowledge graph-supported intelligent lecturing assistant (ILA) system designed to help teachers enhance student learning during lectures by integrating insights from cognitive science, neuroscience, and established pedagogical best practices. By leveraging the power of AI, this system aims to empower instructors and create more interactive and engaging learning environments, ultimately contributing to improved student learning outcomes.
@@ 1 Introduction
Moreover, Oakley et al in [2] introduce “Learn it, link it”, another fundamental concept of learning grounded in neuroscience research. The concept explains the establishment, reinforcement, and expansion of neural connections in the neocortex of the learner’s brain. In educational contexts, two critical types of memory are working memory, which temporarily holds information, and long-term memory, where information can potentially persist for a lifetime. A primary goal of teaching is to facilitate the transfer of content from working memory to long-term memory. Given the limited capacity of working memory, effective teaching method requires breaking down and reviewing content at optimally short intervals to ensure its successful integration and retention in long-term memory.
Figure 1: Overview of Intelligent Lecturing Assistant (ILA) System (The Cartoon Images Generated by MS Copilot.) A knowledge graph that compiles course content and effective teaching strategies forms the system’s foundation. The system constantly analyzes input from both the lecturer and students, including the lecturer’s voice for sentiment, the actual content being delivered, and student responses through various means. In the context of teaching and learning, the system applies the analysis results to assist the teacher to implement the evidence-based learning strategies like retrieval practice, spaced practice, interleaving, and feedback-driven metacognition.
@@
Each of these scientifically grounded strategies and methods involves deliberate practice and reflection, with the goal of helping students in developing persistent long-term memory and critical thinking skills. To implement these strategies and methods effectively, specific techniques and applications are required. In particular, a teacher needs to design low-stakes quizzes, formulate content-relevant practice tests, establish study schedules that revisit material at progressively increasing intervals, split learning into shorter and more frequent sessions, and offer timely and specific feedback on quizzes and assignments. Developing an intelligent lecturing assistant system to support educators in applying these teaching strategies is both practical and beneficial.
@@
Similar to intelligent tutorial systems (ITS), an intelligent lecturing assistant (ILA) also aims to substantially enhance the educational experience. Differently, an ILA will focus on assisting teachers with timely suggestions for pauses and relevant quizzes and questions to reinforce student learning and memory transfer. However, many challenges remain. One of challenges we investigate here is how to balance the human and computer decision-making [17]. To tackle this, the intelligent agent requires comprehensive knowledge of the curriculum, the capability to detect speech sentiment, the ability to compare delivered content with the planned curriculum in real time, the capacity to predict student engagement levels, and the ability to dynamically generate interventions at optimal moments. Knowledge graphs, machine learning, and speech recognition technologies play crucial roles in the development of such a system.
In this paper, we introduce an intelligent lecturing assistant (ILA) system that leverages a knowledge graph to capture course content and enable the reasoning for intervention. Figure 1 illustrates an overview of the system. At its core lies a knowledge graph that compiles course content and effective teaching strategies. The system constantly analyzes input from both the lecturer and students, including the lecturer’s voice for sentiment, the actual content being delivered, and student responses through various means. In the context of teaching and learning, the system applies the analysis results to assist the teacher to implement the evidence-based learning strategies like retrieval practice, spaced practice, interleaving, and feedback-driven metacognition.
We will leverage modern AI techniques including large language models (LLMs), to construct the knowledge graph on the fly from uploaded course content and generate timely interventions based on detected moments. As an initial investigation in this long-term research project, this paper presents two primary areas of content. First, we outline the process of constructing a knowledge graph that provides a structured representation of the content. We explain how this integrates with the representation of lecture sentiment and pedagogical methods.
Second, we detail the development of a predictive model that can classify lecture voice sentiment as engaging or not engaging. we developed a training set comprising over 3,000 one-minute lecture voice clips. Each clip was manually labeled as either engaging or non-engaging. Utilizing this dataset, we constructed and evaluated several classification models based on a variety of features extracted from the voice clips. The results demonstrate promising performance, achieving an F1-score of 90% for boring lectures on an independent set of over 800 test voice clips. This case study lays the groundwork for the development of a more sophisticated model that will integrate content analysis and pedagogical practices. Our ultimate goal is to aid instructors in teaching more engagingly and effectively by leveraging modern artificial intelligence techniques.
The rest of the paper is organized as follows. Section 2 discusses related work. Section 3 describes the process of constructing the knowledge graph for the intelligent lecturing system. Section 4 formalizes the problem of lecture voice sentiment analysis. Section 5 details the process of collecting training and independent validation data. Section 6 discusses the features used to label lecture clips. Section 7 describes various classification models. Section 8 presents the evaluation results. Section 9 discusses the findings and makes suggestions. Section 10 concludes the paper. Finally, Section 11 presents the link to the public Github repository containing the source code and the method for downloading the training and validation data sets.
2 Related Work
Knowledge graph (KG) represents the knowledge in a domain using a graphical structure consisting of (typed-) vertices and (typed-) links [18, 19, 20]. They have gained significant attention for their ability to represent complex information and enhance various AI applications. Knowledge graphs are pivotal in organizing and conveying vast amounts of data, making them invaluable in fields such as education, scientific research, social networks, and more [21].
Knowledge graphs have been increasingly applied in educational contexts to improve both face-to-face and online learning experiences [22, 23, 24]. They facilitate the organization and retrieval of educational content, enhance personalized learning, and support intelligent tutoring systems. By structuring educational data, knowledge graphs enable the development of adaptive learning systems that can tailor content to individual learners’ needs. The review in [22] highlights the role of KGs in personalized learning, curriculum design, concept mapping, and educational content recommendation systems. The research in [23] discusses the fundamental research methods in knowledge maps and analyzes classic cases in the education field. The study in [24] compares different KG models utilized in the education domain.
KGs can be constructed manually or automatically. The studies in [25, 26] discuss the automatic construction of educational KGs. Recently, a great deal of effort has been put into applying pre-trained large language models (LLMs) for constructing knowledge graphs (KGs) [27, 28]. The research in [29] reviews methods for generative KG construction. The studies in [30, 31] explore the (semi-)automatic construction of KGs facilitated by open-source LLMs. Applying LLMs to automatically construct KGs for course content and teaching delivery remains under-explored.
KGs are widely used for information retrieval [32] and question answering [33, 34]. An intelligent lecturing assistant system involves multimodal information including KGs, voice, text, and images. KGs play an important role in retrieving and reasoning relevant information for lecture assistant. A latest study has revealed that generative AI/LLM-driven context personalization positively affected learning motivation [35]. A critical aspect is to combine the strengths of KGs and LLMs for accurate intervention [36].
Intelligent educational systems (IES) and intelligent tutoring systems (ITS) have leveraged KGs to address several key aspects of learning, including exercise recommendation and selection [37], student interactions diagnosis [38], personalized learning guidance [39], multi-dimensional knowledge representation [37], integration with other AI techniques [37, 38], and automatic grading [39]. By leveraging the structured representation of knowledge and the relationships between concepts, these systems can better assess student knowledge, recommend appropriate learning materials, and guide the learning process.
@@ 2 Related Work
To assist teachers in real-time lecturing, an intelligent lecturing assistant system must integrate and reason about the multimodal information available in the classroom. One crucial aspect of this multimodal data is the sentiment conveyed through the lecturer’s voice. The integration of voice sentiment analysis in educational technology builds upon foundational research in affective computing and speech processing [40, 41, 42]. Previous studies have demonstrated that emotional cues in speech, such as tone, pitch, and rhythm, can significantly impact communication efficacy and audience engagement [43]. Speech emotion recognition (SER) is a task to determine the speaker’s emotional state, such as happiness, anger, sadness, or frustration, from speech patterns like tone, pitch, and rhythm [44]. Deep learning approaches have become prominent solutions for SER in recent years. Various neural network architectures have been applied [45]. Feature extraction is a crucial step in SER. Common features used include Mel-frequency Cepstral Coefficients (MFCCs) and spectral features [46, 47]
In educational settings, a teacher’s emotional expressiveness has been linked to student motivation, comprehension, and retention of information [48]. Hence, accurately detecting and responding to these vocal sentiments can be pivotal in creating a more engaging lecturing environment. In this study, we focus on a binary lecture voice sentiment analysis as engaging or not engaging. To build an effective machine learning model, we need a large set of training and test data for benchmarking. We will build such a dataset as a benchmark for future researchers to investigate and improve upon. This dataset is publicly available upon request [^1] https://github.com/anyuanay/KG_ILA to facilitate further advancements in the field of lecture sentiment analysis and intelligent lecturing assistant systems.
3 Constructing Knowledge Graph for Intelligent Lecturing System
The knowledge graph, which represents the course content, schedule, objectives, and assessment is the foundation of the proposed intelligent lecturing assistant (Figure 1). The key requirements for the construction and representation of this knowledge graph (KG) include:
  • Accurate Representation: The KG must accurately capture the detailed content, schedule, objectives, and assessment materials of the course.
  • Automatic Construction: The KG should be automatically constructed for a wide range of topics across various educational contexts.
  • Support for Effective Teaching: The KG should support to the generation of relevant quizzes and questions to engage students at optimal time.
  • Support for Improving Learning: The KG should enable reasoning with pedagogical strategies to enhance teaching effectiveness and support student learning.
Prior to constructing the knowledge graph, we need to develop an ontology to provide a structured framework and common understanding of the domain. An ontology defines the key concepts and their relationships that guide the creation of the knowledge graph. The development of ontology is essential for the intelligent lecturing assistant, as it enables the efficient management, organization, and utilization of educational content. Existing ontology [49] models university curricula and students’ profiles. It focuses on curriculum recommendation for college students. For lecturing assistant, we propose a lecture ontology presented in Figure 2.
Figure 2: A Lecture Ontology for Intelligent Lecturing Assistant
The lecture ontology encompasses primary concepts such as course, course module, schedule, and assessment, along with their subclasses, including section, lecture, topics, assignment, quiz, test, and project. The design of this ontology aims to facilitate the implementation of the evidence-based strategies including retrieval practices, spaced practice, interleaving, and feedback-driven metacognition.
With the ontology in place, we will apply pre-trained large language models (LLM) to extract instances from course materials to construct the knowledge graph. The extraction procedure begins by processing the course materials, which may include lecture notes, textbooks, syllabi, and other instructional documents. The procedure scans the text, identifying instances of the defined concepts and their interrelationships. For example, it may extract information about a course module, including its sections and topics, and link it to the appropriate schedule. Similarly, it can identify assessment types and schedules, mapping them to the relevant course modules. Once the instances are extracted, they are integrated into the knowledge graph according to the ontology’s structure. This integration involves linking the extracted instances to their corresponding concepts and establishing relationships as defined by the ontology. For instance, a "lecture" instance would be linked to its parent "course module," and an "assessment" instance would be associated with the relevant "course module" and "objective."
Given the knowledge graph as the foundation, the intelligent lecturing assistant will listen to the lecture and determine the optimal moments to assist the instructor in implementing appropriate educational strategies. The system will employ the following four rules to guide its interventions:
  • Rule 1 - applying retrieval practice: The assistant will generate low-stake quizzes from topics covered in the previous two weeks. These quizzes will prompt students to recall information, reinforcing their memory and understanding of recent material.
  • Rule 2 - applying spaced practice: The assistant will create mixed quiz questions from the previous four weeks, focusing on similar topics. This approach helps reinforce learning by revisiting material at intervals, promoting long-term retention.
  • Rule 3 - applying interleaving: The assistant will incorporate questions from different topics from the previous four week within the same quiz. This method challenges students to distinguish between concepts and apply their knowledge flexibly, improving problem-solving skills and transfer of learning.
  • Rule 4 - applying feedback-driven metacognition: The assistant will guide students to identify what they know and what they don’t know. The intervention will highlight areas of strength and weakness, encouraging students to adjust their study strategies and enhancing their metacognition skills.
To apply these rules effectively, the intelligent lecturing assistant (ILA) must detect the optimal moment by analyzing the sentiment in the lecture speech. In the rest of the paper, we describe the process of training a machine learning model capable of classifying the lecture voice sentiment as engaging or not engaging.
4 The Problem of Lecture Voice Sentiment Analysis
@@
The problem of lecture voice sentiment analysis is to classify whether the lecture speech in a certain period of time as engaging or not engaging. Formally, Let $S$ represent a lecture speech with a duration of $T$ minutes. Let $y$ denote the label associated with speech $S$, where $y\in\{0,1\}$. Here, $y=0$ indicates that the lecture is "engaging," and $y=1$ indicates that the lecture is "not engaging." We want to train a model $M$ such as $M(S)=y\in\{0,1\}$.
@@
5 Building Data Sets for Lecture Voice Sentiment Analysis
5.1 Training Data
@@
Finally, we collected total 3,025 lecture voice clips. Table 1 shows the numbers of clips collected by each researcher. Figure 3 illustrates the distribution of the clips labeled as ‘engaging’ and ‘boring’ by the researchers. It should be noted that the target labels are evenly distributed in the collected data set for the binary classification problem.
5.2 Independent Validation Data
@@
Researcher Data Collection Dates Count of Engaging Clips Count of Boring Clips
No.1 06/17/2024-07/12/2024 644 363
No.2 06/17/2024-07/12/2024 201 233
No.3 06/17/2024-07/12/2024 185 534
No.4 06/17/2024-07/12/2024 200 200
No.5 06/17/2024-07/12/2024 217 248
Table 1: Data Counts Collected by Researchers. We had in total 5 researchers (No.1 - 5) who contributed to the data collection process.
@@
6 Extracting Features from Raw Voice Signals
@@
The raw audio signal is typically divided into small, overlapping frames to analyze the time-varying spectral properties. For a feature that is computed across frames, we compute the mean across all frames to obtain a more stable representation for the entire audio clip. Table 2 shows the feature names and their final dimensions.
@@
Table 2: Features and Dimensions; the Total Dimension is 209
@@
7 Building Models for Lecture Voice Sentiment Analysis
@@
For deep neural networks, we switch between a fully connected neural network and a convolutional neural network depending the dimension of the input features. If the input dimension is smaller than 20, we build a fully connected neural networks where multiple dense layers are stacked to create a deep network. Each layer captures different levels of abstraction, allowing the model to learn complex representations from the input features. If the input dimension is at least 20, we experiment with a convolution neural network whose architecture can capture local patterns in the feature space, which might be crucial for distinguishing between engaging and non-engaging lecture clips.
Table 3 show the summary of the models we explored in the study.
@@
Table 3: The Summary of the Traditional Models and the Neural Networks Evaluated for Lecture Voice Sentiment Analysis
8 Experimental Results

### v2

# Is the Lecture Engaging? Lecture Sentiment Analysis for Knowledge Graph-Supported Intelligent Lecturing Assistant (ILA) System

@@ Abstract
This paper introduces an intelligent lecturing assistant (ILA) system that utilizes a knowledge graph to represent course content and optimal pedagogical strategies. The system is designed to support instructors in enhancing student learning through real-time analysis of voice, content, and teaching methods. As an initial investigation, we present a case study on lecture voice sentiment analysis, in which we developed a training set comprising over 3,000 1-minute lecture voice clips. Each clip was manually labeled as either engaging or non-engaging. Utilizing this dataset, we constructed and evaluated several classification models based on a variety of features extracted from the voice clips. The results demonstrate promising performance, achieving an F1-score of 90% for boring lectures on an independent set of over 800 test voice clips. This case study lays the groundwork for the development of a more sophisticated model that will integrate content analysis and pedagogical practices. Our ultimate goal is to aid instructors in teaching more engagingly and effectively by leveraging modern artificial intelligence and big data techniques.
§§footnotetext: Drexel summer interns contributing equally to this project.
I Introduction
Engaging students and facilitating the retention of knowledge in long-term memory are fundamental aspects of effective teaching, yet many lectures often fail to achieve this goal. Although extensive research in cognitive science [1] and neuroscience [2] has suggested a multitude of scientifically-based strategies for effective teaching, the application of these findings in real classrooms remains limited. Advancements in artificial intelligence and big data offer a promising avenue for transforming the science of learning to real-world experience. This paper introduces a novel knowledge graph-supported intelligent lecturing assistant (ILA) system designed to help teachers enhance student learning during lectures by integrating insights from scientifically-based teaching strageties. By leveraging the power of AI, this system aims to empower instructors and create more interactive and engaging learning environments, ultimately contributing to improved student learning outcomes.
@@ I Introduction
Additionally, Oakley et al in [2] introduce “Learn it, link it”, another fundamental concept of learning grounded in neuroscience research. The concept explains the establishment, reinforcement, and expansion of neural connections in the neocortex of the learner’s brain. In educational contexts, learning involves two critical types of memory: working memory, which temporarily holds information, and long-term memory, where information can potentially persist for a lifetime. A primary goal of teaching is to facilitate the transfer of content from working memory to long-term memory. Given the limited capacity of working memory, effective teaching method requires breaking down and reviewing content at optimally short intervals to ensure its successful integration and retention in long-term memory.
@@
Each of these scientifically grounded strategies and methods involves deliberate practice and skills. Specifically, a teacher needs to design low-stakes quizzes, formulate content-relevant practice tests, establish study schedules that revisit material at progressively increasing intervals, split learning into shorter and more frequent sessions, and offer timely and specific feedback on quizzes and assignments. Developing an intelligent lecturing assistant system to support teachers in applying these teaching strategies is both practical and beneficial. However, for such a system to be effective, it must be capable of analyzing lecture sentiment and determining the optimal moments for retrieval practice. Additionally, the system must possess comprehensive knowledge of the entire curriculum to effectively apply spacing and interleaving techniques.
In this paper, we detail the development of an approach that can classify lecture voice sentiment for an intelligent lecturing assistant (ILA) system which leverages a knowledge graph to represent the curriculum and learning objectives. With the insights from the lecture sentiment analysis, the ILA system will be able to assist teachers with timely suggestions for pauses and relevant quizzes and questions to reinforce student learning and memory transfer. Our main contributions in this study include:
  • We created a unique training data set comprising over 3,000 1-minute lecture voice clips, and a unique validation set of over 800 independent voice clips. Each clip was manually labeled as either engaging or non-engaging for sentiment analysis.
  • We extracted a range of features from the voice clips, including temporal characteristics, perceptual representations, spectral features, and pitch and harmonic content.
  • Utilizing this dataset and features, we constructed and evaluated a variety of classification models. The results demonstrate promising performance, achieving an F1-score of 90% for boring lectures on the independent validation set.
  • This case study lays the groundwork for the development of a more sophisticated model that will integrate content analysis and pedagogical practices. Our ultimate goal is to aid instructors in teaching more engagingly and effectively by leveraging modern artificial intelligence and big data techniques.
The rest of the paper is organized as follows. Section II discusses related work. Section III presents a background for the intelligent lecturing system. Section IV formalizes the problem of lecture voice sentiment analysis. Section V details the process of collecting training and independent validation data. Section VI presents a range of features extracted from raw acoustic signals for down-stream classification. Section VII describes various classification models. Section VIII presents the evaluation results. Section IX discusses the findings and makes suggestions. Section X concludes the paper. Finally, the Appendix presents the link to the public Github repository containing the source code and the method for downloading the training and validation data sets.
II Related Work
@@
In educational settings, a teacher’s emotional expressiveness has been linked to student motivation, comprehension, and retention of information [17]. Hence, accurately detecting and responding to these vocal sentiments can be pivotal in creating a more engaging lecturing environment. The integration of voice sentiment analysis in educational technology builds upon foundational research in affective computing and speech processing [18, 19, 20]. Previous studies have demonstrated that emotional cues in speech, such as tone, pitch, and rhythm, can significantly impact communication efficacy and audience engagement [21]. Speech emotion recognition (SER) is a task to determine the speaker’s emotional state, such as happiness, anger, sadness, or frustration, from speech patterns like tone, pitch, and rhythm [22]. Deep learning approaches have become prominent solutions for SER in recent years. Various neural network architectures have been applied [23]. Feature extraction is a crucial step in SER. Common features used include Mel-frequency Cepstral Coefficients (MFCCs) and spectral features [24, 25].
III Background
Figure 1: Overview of Intelligent Lecturing Assistant (ILA) System. A knowledge graph that compiles course content, schedule, and learning objectives forms the system’s foundation. The system constantly analyzes input from both the lecturer and students, including the lecturer’s voice for sentiment, the actual content being delivered, and student responses through various means. In the context of teaching and learning, the system applies the analysis results to assist the teacher to implement the evidence-based learning strategies like retrieval practice, spaced practice, interleaving, and feedback-driven metacognition.
In this section, we provide an overview of the knowledge graph-supported intelligent lecturing assistant (ILA) system, with lecture sentiment analysis functioning as a critical component. Figure 1 illustrates the overview of the system. At its core lies a knowledge graph that compiles course content, schedule, and learning objectives. The system constantly analyzes input from both the lecturer and students, including the lecturer’s voice for sentiment, the actual content being delivered, and student responses through various means. In the context of teaching and learning, the system applies the analysis results to assist the teacher to implement the scientifically-based teaching strategies like retrieval practice, spaced practice, interleaving, and feedback-driven metacognition. Knowledge graphs, machine learning, and speech recognition technologies play crucial roles in the development of such a system.
Intelligent educational systems (IES) and intelligent tutoring systems (ITS) have leveraged knowledge graphs to address several key aspects of learning, including exercise recommendation and selection [26], student interactions diagnosis [27], personalized learning guidance [28], multi-dimensional knowledge representation [26], integration with other AI techniques [26, 27], and automatic grading [28]. By leveraging the structured representation of knowledge graphs, these systems can better assess student knowledge, recommend appropriate learning materials, and guide the learning process.
IV The Problem of Lecture Voice Sentiment Analysis
@@
The problem of lecture voice sentiment analysis is to classify whether the lecture speech in a certain period of time as engaging or not engaging. Formally, Let $S$ represent a lecture speech with a duration of $T$ minutes. Let $y$ denote the label associated with speech $S$, where $y\in\{0,1\}$. Here, $y=0$ indicates that the lecture is ”engaging,” and $y=1$ indicates that the lecture is ”not engaging.” We want to train a model $M$ such as $M(S)=y\in\{0,1\}$.
@@
V Building Data Sets for Lecture Voice Sentiment Analysis
V-A Training Data
@@
Finally, we collected total 3,025 lecture voice clips. Table I shows the numbers of clips collected by each researcher. Figure 2 illustrates the distribution of the clips labeled as ‘engaging’ and ‘boring’ by the researchers. It should be noted that the target labels are evenly distributed in the collected data set for the binary classification problem.
V-B Independent Validation Data
@@
Researcher Collection # Engaging Clips # Boring Clips
Date
No.1 06/17/2024- 644 363
07/12/2024
No.2 06/17/2024- 201 233
07/12/2024
No.3 06/17/2024- 185 534
07/12/2024
No.4 06/17/2024- 200 200
07/12/2024
No.5 06/17/2024- 217 248
07/12/2024
TABLE I: Data Counts Collected by Researchers. We had in total 5 researchers (No.1 - 5) who contributed to the data collection process.
@@
VI Extracting Features from Raw Voice Signals
@@
The raw audio signal is typically divided into small, overlapping frames to analyze the time-varying spectral properties. For a feature that is computed across frames, we compute the mean across all frames to obtain a more stable representation for the entire audio clip. Table II shows the feature names and their final dimensions.
@@
TABLE II: Features and Dimensions; the Total Dimension is 209
@@
VII Building Models for Lecture Voice Sentiment Analysis
@@
For deep neural networks, we switch between a fully connected neural network and a convolutional neural network depending the dimension of the input features. If the input dimension is smaller than 20, we build a fully connected neural networks where multiple dense layers are stacked to create a deep network. Each layer captures different levels of abstraction, allowing the model to learn complex representations from the input features. If the input dimension is at least 20, we experiment with a convolution neural network whose architecture can capture local patterns in the feature space, which might be crucial for distinguishing between engaging and non-engaging lecture clips. We selected 20 as the minimum number of dimensions for switching between a fully connected neural network and a convolutional neural network because of empirical considerations and avoiding overfitting.
Table III show the summary of the models we explored in the study.
@@
TABLE III: The Summary of the Traditional Models and the Neural Networks Evaluated for Lecture Voice Sentiment Analysis
VIII Experimental Results